The following table of values gives a company's annual profits in millions of dollars. Rescale the data so that the year 2003 corresponds to x = 0.

Year 2003 2004 2005 2006 2007 2008
Profits (in millions of dollars) 51.7 53.2 51.9 52.2 56.1 58.3


Find the quadratic regression model for this data.

a) y=0.45x2−0.19x+51.98

b) y=0.4339x2−0.9696x+52.3464

c) y=0.4339x2−1.8375x+53.75

d) y=0.7571x2−1.5886x+52.9743

e) y=0.4571x2−1.0486x+52.3743

f) None of the above
Answer #1

from given data:

y x x^2
51.7 0 0
53.2 1 1
51.9 2 4
52.2 3 9
56.1 4 16
58.3 5 25

Applying multiple regression from data analysis tab of excel :

below is regression equation:

y^ =0.4339*X2-0.9696X+52.3464

option B is correct
Answer #2

Answer with Calculation:

  1. Rescale Years to x=0 starting at 2003:
    x-values: 0,1,2,3,4,5.

  2. Test Key Points for Each Option:

    • Option a:
      x=0y=51.98 (vs. 51.7) → Error: +0.28
      x=1y=52.24 (vs. 53.2) → Error: -0.96
      Poor fit.

    • Option b:
      x=0y=52.3464 (vs. 51.7) → Error: +0.65
      Immediate mismatch.

    • Option c:
      x=0y=53.75 (vs. 51.7) → Error: +2.05
      Invalid.

    • Option d:
      x=0y=52.9743 (vs. 51.7) → Error: +1.27
      Poor start.

    • Option e:
      x=0y=52.3743 (vs. 51.7) → Error: +0.67
      x=1y=51.78 (vs. 53.2) → Error: -1.42
      Does not align.

  3. Conclusion:
    All options show significant errors (>0.5) at critical points (e.g., x=0 and x=1).

Final Answer:
                               f). None of the above 


source: Self
answered by: Harshwardhan kunal